OpenStudy (anonymous):

A simple question to kill the time#11 find 3+2+1+0.11+0.0011+0.000011=? a.frac{45}{7} b.frac{55}{9} c.frac{67}{10} d.frac{79}{13}

OpenStudy (anonymous):

a.45/7 b.55/9 c.67/10 d.79/13

OpenStudy (anonymous):

whatever 6 + 1/9 is

OpenStudy (anonymous):

yea easy.. use sum up to inifinity.. :D

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

i guess that would be b

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

sum shmum just recalled that \[\overline{.1}=\frac{1}{9}\]

OpenStudy (anonymous):

for an elaborate solution its 6+ 11( 1/100 + 1/10000 + 1/1000000 and so on) the term inside bracket is a GP with a = 1/100 and r= 1/100.. so find sum upto infinity!